# AI-Assisted Transport of Radioactive Ion Beams

### v2

@@ I Introduction
In this work, we introduce a system that uses Artificial Intelligence (AI) to assist in the radioactive beam transport process. We present results from applying our methodology to real-life scenarios where the observable quantity to maximize is the activity from beta decay of stopped radioactive ions and where electrostatic devices are used to transport the beam. We also discuss the results and advantages when compared with standard tuning methods, and suggest that our method could be applied to similar radioactive beam facilities worlwide.
@@ III Methodology
Given these technical challenges, the use AI applications, and in particular Machine Learning (ML) methods [11, 12], hold the potential to achieve complete automation of beam extraction, delivery, and optimization of related instruments. The aim of the work presented here is to automate radioactive beam tuning using ML techniques. Specifically, the use of Bayesian Optimization (BO) for control and tuning of beam elements. Although the underlying theory of this optimization approach is not new [13], computational advances and open-source software libraries [14, 15, 16] have only recently matured enough to consider practical applications.
@@
Early successes applying ML in the form of BO [18, 19, 20, 21, 22, 23, 24] has proven the effectiveness of this method. This technique has been chosen over others such as Reinforcement Learning (RL) [25] due to its non-parametric nature. GP does not require pre-training a Neural Network (NN) on either simulated or historical data. The learning is in real time from current samples and the model estimations improve with each iteration. Although digital-twins of the system could be useful as a testbed for new algorithms, they fail to account for the underlying changing dynamics and fringe behaviors. Likewise, historical data is unable to accurately represent current parameters. Thus, BO is the most straightforward method to automate a complex system – like the transport of radioactive beams – without a priori knowledge.
@@ IV Implementation
As described in Ref. [26], variables, variable ranges, and the observables are specified in an abstract element called Environment. Variables, such as DC voltages controlling steerers, quadrupoles, and focusing elements, are set within operational limits. Historical “good-tune” values are loaded at initialization. The observable, defined as the optimization target (labeled “Feedback” in Fig. 2), uses signals from particle detectors. In the case of RIBs produced by CARIBU, we used Silicon detectors to observe electrons from the beta-decay of the isotope of interest. Three high-level class methods facilitate communication with devices: reading current values, setting new values from optimization, and obtaining feedback for optimization. A second abstract element, called Interface, contains the instructions for device communication.
@@ V Results
Figure 3 illustrates various aspects of optimizing the radioactive beamline Section 1. The top panel displays live optimization of the observed activity versus iteration. The times scale of our optimization iterations is selected such that, after background subtraction, an increment of the observed activity is proportional to an increment in the RIB intensity. Initially, the observed rate fluctuates as the algorithm explores the tuning space of quadrupoles and steerers, characterized by large changes. As iterations progress, the algorithm refines the machine configuration, resulting in steady RIB rate improvements. Despite significant drops around iterations 70, 80, and 100, the algorithm successfully returned to optimal settings. In the final iteration, a background measurement was conducted by removing the beam to measure the beam-off rate, stemming from the beta decay chain. The Bateman equation [30] for a two-isotope chain approximates this curve:
Figure 3: Live optimization of radioactive ion transport in beamline Section 1. In the top panel, the activity in counts per second (cps) is shown as a function of the optimization iteration. A yellow star marks the maximum, which corresponds to the optimum configuration. The background from beta decay is shown in light blue. The bottom panel shows variable changes corresponding to the top panel activity. Each iteration took about 10 seconds.
$\frac{\lambda_{P}}{\lambda_{D}-\lambda_{P}}A_{P}(0)(e^{-\lambda_{P}t}-e^{-\lambda_{D}t})+A_{D}(0)e^{\lambda_{D}t}$ (6)
Where $\lambda_{P}$ is the decay constant of the parent nucleus, $\lambda_{D}$, the decay constant of the daughter nucleus, $A_{P}$ is the initial beam rate, and $A_{D}$ is the initial daughter rate. The net rate is calculated as: (Observed Activity) - (Background Activity).
@@
Strong positive correlations (red) indicate direct proportionality, while negative correlations (blue) suggest inverse relationships. For instance, QH-control and Steer1-X have small positive correlations (0.336 and 0.280) with Beam Rate, indicating moderate increases with voltage. Conversely, Steer1-Y, Quad1-X, and Quad1-Y show high negative correlations (-0.706, -0.647, -0.547), suggesting beam rate benefits from reduced voltage. A high positive correlation (0.891) between Beam Rate and Iteration aligns with optimization rules. Elements like Steer2-X (-0.003), Steer2-Y (0.024), and QS-Y (-0.150) show near-zero correlation, indicating no direct influence on beam rate.
@@
It is important to note that while ARD captures potentially nonlinear and interaction-based effects within the model, the correlation matrix reflects empirical, linear relationships within the observed data. Due to the non-uniform sampling nature of Bayesian Optimization—where the search is biased toward promising regions—the correlation values may not directly correspond to global sensitivities. However, strong correlations often indicate variables that contributed significantly to observed changes in beam rate, and can serve as a complementary tool to the ARD-learned lengthscales. Further analysis identifies Quad1-X (-0.647), Quad1-Y (-0.547), Steer1-Y (-0.706), and QH-control (0.336) as the most sensitive elements to beam rate. By focusing on these key elements, we can enhance optimization efficiency by reducing unnecessary variables, ultimately simplifying the troubleshooting process.
@@
Figure 5: Effect of varying $\beta$ on optimization. Left: Conservative exploration with $\beta=2$ stalls at a suboptimal local maximum (observed no RIB). Center: Aggressive exploration with $\beta=3$ reaches high-performance regions (found RIB). Right: Returning to $\beta=2$ prioritizes exploitation, converging to the global maximum (increased RIB rate).
Initially, a conservative approach with $\beta=2$ resulted in the algorithm stalling at a suboptimal local maximum, as no RIB was observed. By increasing $\beta$ to 3, the algorithm adopted a more aggressive exploration strategy, successfully identifying high-performance regions and detecting the RIB. Subsequently, reverting to $\beta=2$ allowed the algorithm to focus on exploitation, converging to the global maximum and achieving an increased RIB rate.

### v3

@@ I Introduction
In recent years, there has been increasing interest in using Artificial Intelligence (AI) applications and Machine Learning (ML) methods to improve the automation of accelerator operations. A review paper by Roussel et al. [11] highlights how Bayesian Optimization algorithms offer an effective and adaptable approach to addressing a wide range of optimization challenges in accelerator physics.
In this work, we introduce a system that uses AI to assist in the transport of radioactive ion beams. We present results from applying our methodology to real-life scenarios where the observable quantity to maximize is the activity from beta decay of stopped radioactive ions and where electrostatic devices are used to transport the beam. We also discuss the results and advantages when compared with standard tuning methods, and suggest that our method could be applied to similar radioactive beam facilities worldwide.
@@ III Methodology
Given these technical challenges, the use AI applications, and in particular ML methods [12, 13], hold the potential to achieve complete automation of beam extraction, delivery, and optimization of related instruments. The aim of the work presented here is to automate radioactive beam tuning using ML techniques. Specifically, the use of Bayesian Optimization (BO) for control and tuning of beam elements [11]. Although the underlying theory of this optimization approach is not new [14], computational advances and open-source software libraries [15, 16, 17] have only recently matured enough to consider practical applications.
@@
Early successes in applying ML in the form of BO [19, 20, 21, 22, 23, 24, 25] have proven the effectiveness of this method. This technique has been chosen over others such as Reinforcement Learning (RL) due to its non-parametric nature and minimal prior data requirement [26, 27]. GP does not require pre-training a Neural Network (NN) on either simulated or historical data. The learning is in real time from current samples, and the model estimations improve with each iteration. Although digital-twins of the system could be useful as a testbed for new algorithms, they fail to account for the underlying changing dynamics and fringe behaviors. Likewise, historical data is unable to accurately represent current parameters. Thus, BO is the most straightforward method to automate a complex, dynamic system – like the transport of radioactive beams – without a priori knowledge.
@@ IV Implementation
As described in Ref. [28], variables, variable ranges, and the observables are specified in an abstract element called Environment. Variables, such as DC voltages controlling steerers, quadrupoles, and focusing elements, are set within operational limits. Historical “good-tune” values are loaded at initialization. The observable, defined as the optimization target (labeled “Feedback” in Fig. 2), uses signals from particle detectors. In the case of RIBs produced by CARIBU, we used Silicon detectors to observe electrons from the beta decay of the isotope of interest. Three high-level class methods facilitate communication with devices: reading current values, setting new values from optimization, and obtaining feedback for optimization. A second abstract element, called Interface, contains the instructions for device communication.
@@ V Results
Figure 3 illustrates various aspects of optimizing the radioactive beamline Section 1. The top panel displays live optimization of the observed activity versus iteration. The times scale of our optimization iterations is selected such that, after background subtraction, an increment of the observed activity is proportional to an increment in the RIB intensity. Initially, the observed rate fluctuates as the algorithm explores the tuning space of quadrupoles and steerers, characterized by large changes. As iterations progress, the algorithm refines the machine configuration, resulting in steady RIB rate improvements. Despite significant drops around iterations 70, 80, and 100, the algorithm successfully returned to optimal settings. In the final iteration, a background measurement was conducted by removing the beam to measure the beam-off rate, stemming from the beta decay chain of the isotopes involved. This background is crucial for accurately determining the net activity of the beam.
The Bateman equation [32] is a mathematical model used to describe the decay of isotopes in a sequential chain, where a parent isotope decays into a daughter isotope, which may further decay into subsequent isotopes. In our case, the equation is applied to a two-isotope chain, capturing the decay dynamics of the parent and daughter isotopes involved in the beta decay process.
The equation is given by:
Figure 3: Live optimization of radioactive ion transport in beamline Section 1. In the top panel, the activity in counts per second (cps) is shown as a function of the optimization iteration. A yellow star marks the maximum, which corresponds to the optimum configuration. The background from beta decay, modeled using the Bateman equation, is shown in light blue. The bottom panel shows variable changes corresponding to the top panel activity. Each iteration took about 10 seconds.
$A_{B}=\frac{\lambda_{P}}{\lambda_{D}-\lambda_{P}}A_{P}(0)(e^{-\lambda_{P}t}-e^{-\lambda_{D}t})+A_{D}(0)e^{\lambda_{D}t}$ (6)
where $A_{B}$ is the background activity, $\lambda_{P}$ is the decay constant of the parent isotope, $\lambda_{D}$, the decay constant of the daughter isotope, $A_{P}$ is the initial beam rate, and $A_{D}$ is the initial daughter isotope rate. The Bateman equation allows us to predict the activity levels of these isotopes over time, providing a theoretical curve that represents the background activity due to beta decay. This curve, shown in light blue in the top panel of Fig. 3, is subtracted from the observed activity to calculate the net rate: (Observed Activity) - (Background Activity). By modeling the background activity accurately, we can isolate the true signal from the background.
@@
Strong positive correlations (red) indicate direct proportionality, while negative correlations (blue) suggest inverse relationships. For instance, QH-control and Steer1-X have small positive correlations (0.336 and 0.280) with Beam Rate, indicating moderate increases with voltage. Conversely, Steer1-Y, Quad1-X, and Quad1-Y show high negative correlations (-0.706, -0.647, -0.547), suggesting beam rate benefits from reduced voltage. Elements like Steer2-X (-0.003), Steer2-Y (0.024), and QS-Y (-0.150) show near-zero correlation, indicating no direct influence on beam rate.
@@
It is important to note that while ARD captures potentially nonlinear and interaction-based effects within the model, the correlation matrix reflects empirical, linear relationships within the observed data. Due to the non-uniform sampling nature of Bayesian Optimization—where the search is biased toward promising regions—the correlation values may not directly correspond to global sensitivities. However, strong correlations often indicate variables that contributed significantly to observed changes in beam rate, and can serve as a complementary tool to the ARD-learned lengthscales.
While our correlation analysis provides empirical insights into linear relationships between tuning variables and beam rate, it is important to consider how this approach relates to existing methodologies. Hanuka et al. [33] explore the use of a physics-informed GP model for optimizing particle accelerators, embedding known physical correlations into the GP framework. This method enhances predictive capabilities by capturing interdependencies between input parameters, informed by physics-based simulations and modeling.
Our work aligns with these principles, utilizing empirical correlations as a complementary tool to ARD. By recognizing the role of correlated kernels, we highlight the potential for integrating physics-informed correlations into our analysis. This suggests avenues for future research to refine optimization strategies further, potentially combining empirical data-driven insights with physics-informed modeling to enhance the robustness of optimization algorithms.
Further analysis identifies Quad1-X (-0.647), Quad1-Y (-0.547), Steer1-Y (-0.706), and QH-control (0.336) as the most sensitive elements to beam rate. By focusing on these key elements, we can enhance optimization efficiency by reducing unnecessary variables, ultimately simplifying the troubleshooting process.
@@
Figure 5: A sequence of events showing how the RIB was found by varying the $\beta$ parameter. Left: Initial case where no RIB is observed and a conservative exploration with $\beta=2$ stalls at a suboptimal local maximum. Center: Since no beam was found after about 100 iterations, a more aggressive exploration with $\beta=3$ was performed, restarting the optimization process from the same initial configuration as the previous run. In less than 20 iterations, the RIB was found, however relatively small gains were made in subsequent iterations. Right: After 80 iterations with $\beta=3$, the optimization was stopped and the parameter $\beta$ was returned to 2, while restarting the process with the variables from the last configuration where the RIB was observed. The optimization was resumed prioritizing exploitation, which resulted in a significant increment in the observed activity (i.e. higher RIB intensity).
Initially, a conservative approach with $\beta=2$ resulted in the algorithm stalling at a suboptimal local maximum, as no RIB was observed. By increasing $\beta$ to 3 and reinitializing the optimization process from a single starting configuration, the algorithm adopted a more aggressive exploration strategy. This change led the algorithm to successfully identify high-performance regions and detect the RIB. Subsequently, reverting to $\beta=2$ and reinitializing the process with the “good-tune” parameters from the $\beta=3$ run, allowed the algorithm to focus on exploitation, converging to the global maximum and achieving an increased RIB rate.